OpenStudy (breathlessreaper):

?help? The functions f(x) and g(x) are described below. f(x) = 2x + 9 g(x) = 2x - 15 The graph of g(x) is obtained by shifting down the graph of f(x) by _____ units.

OpenStudy (the_fizicx99):

How much do you have to shift down f(x) to obtain g(x)? Basically move down from +9 to -15

OpenStudy (anonymous):

@Jesstho.-.

OpenStudy (breathlessreaper):

6?

OpenStudy (anonymous):

Do you want me to explain the solution?

OpenStudy (the_fizicx99):

Erm,  And so on. 9 + 15 = 24, 9 - 24 = -15 so you shift down f(x) = 2x + 9 down 24 intervals.

OpenStudy (anonymous):

Do you still need help @BreathlessReaper

OpenStudy (breathlessreaper):

Yeah I dont really understand o - o

OpenStudy (anonymous):

I'll help :)

OpenStudy (anonymous):

The answer is 24

OpenStudy (anonymous):

And I'll explain it.

OpenStudy (the_fizicx99):

If your are a point that is +9, how many places do you have to move down to get to -15?

OpenStudy (breathlessreaper):

Ohh, wait, okay, im starting to understand @tHe_FiZiCx99 I had to slow down and read it over a few times.

OpenStudy (anonymous):

Since both lines are parallel

OpenStudy (anonymous):

The shifting is equal to difference in y intercept

OpenStudy (the_fizicx99):

A simple way is to subtract them hun, 9 - (-15) = 9 + 15 = 24. So if you subtract 9 - 24 you get -15. You're being told to obtain g(x) using f(x). G(x) is at -15.

OpenStudy (anonymous):

Here y intercept = 9 and -15 thus different = 24

OpenStudy (anonymous):

I hope you understand now :) Have a great evening/night!

OpenStudy (the_fizicx99):

And the answer isn't 24 @Jesstho.-. ... 24 is interpreted at +24. You shift down f(x) -24 intervals to obtain -15 :3

OpenStudy (anonymous):

How'd you get that?

OpenStudy (anonymous):

I got 24.

OpenStudy (anonymous):

@wolfe8

OpenStudy (the_fizicx99):

f(x) = +9 g(x) = -15 You're at 9. You go DOWN to -15, from 9 to 0 it's -9 from 0 to -15 is -15. -9 - 15 = - 24

OpenStudy (the_fizicx99):

http://prntscr.com/3jxqla

OpenStudy (anonymous):

I'm going to redo the problem.

OpenStudy (the_fizicx99):

K have fun. @BreathlessReaper get it? ^-^

OpenStudy (breathlessreaper):

@tHe_FiZiCx99 I believe so. I think i had just made it a lot harder for myself than it should have been xD

OpenStudy (anonymous):

I will :P

OpenStudy (the_fizicx99):

It's really a matter of graphing it and counting down the intervals or just adding them

OpenStudy (anonymous):

Omg,I got 12.

OpenStudy (the_fizicx99):

( ._________________________________________.) @BreathlessReaper there are rules to entering in numerical answer. Some don't allow for signs in them, because it says "down" they're probably implying the answer is already negative (obviously) so putting -24 can be counted incorrect by your exam ._.

OpenStudy (breathlessreaper):

I got it ^^ Thank youuuuuu a billion times c: @tHe_FiZiCx99

OpenStudy (the_fizicx99):

You're welcome :>